OpenStudy (anonymous):

Given a circle with a diameter of 4, what is the circumference?

OpenStudy (mayankdevnani):

do you know what is the formula for finding the circumference of a circle? @nerdybill

OpenStudy (mayankdevnani):

\[\huge circumference~of~circle=\pi \times d\]

OpenStudy (shamim):

diameter \[d=4m\]

OpenStudy (mayankdevnani):



OpenStudy (shamim):

so radius \[r=\frac{ 4 }{ 2 }=2m\]

OpenStudy (shamim):

formula for calculating circumference is\[2 \pi r\]

OpenStudy (shamim):

\[2 \pi \times 2\]

OpenStudy (mayankdevnani):

@shamim use direct method....

OpenStudy (shamim):

ok

OpenStudy (shamim):

hw

OpenStudy (mayankdevnani):

\[\huge \pi \times d~Not~ 2 \times \pi \times r\]